OpenStudy (anonymous):

A rectangle has an area of 180 cm^2 and a perimeter of 54 cm. what are its dimensions?

OpenStudy (wolf1728):

width * height = 180 2w + 2h = 54 w + h = 27 width = 27-h height = 180 / width height = 180 / (27-h) 27*h -h^2 = 180 h = 12 and h = 15 width = 27 -h width = 27 -12 = 15 (when h=12) width = 27 -15 = 12 (when h=15) So, width = 15 height =12 OR width = 13 height =15

OpenStudy (anonymous):

Thanks :)

OpenStudy (wolf1728):

You are welcome :-)

OpenStudy (anonymous):

It's l:15 and w:12

OpenStudy (wolf1728):

Yeah I mistyped my second answer it should be w = 12 h = 15 OR w = 15 h = 12